OpenStudy (anonymous):

how do I graph the equation x-y=5?

OpenStudy (anonymous):

i would begin by writing this equation in terms of y: x-y=5 subtract x from both sides -y=-x+5 now divide both sides by negative 1 y=x-5

OpenStudy (anonymous):

Then i would find the x and y intercepts y=x-5 is linear;)

OpenStudy (anonymous):

ok so how would I find my points? I got that far but I can't find my points to graph or should I just start at the y intercept and guess for another point? given that it is a solutin to the equation?

OpenStudy (anonymous):

nevermind thank you I got it duh my slope is one